46. The pka for acetic acid is 4.76. What is the pH of a 0.04 M solution? (1) (3) 3.03 12.1 5.88 1.40 3.08 (5)
Answer #1

use:

pKa = -log Ka

4.76 = -log Ka

Ka = 1.738*10^-5

CH3COOH dissociates as:

CH3COOH -----> H+ + CH3COO-

4*10^-2 0 0

4*10^-2-x x x

Ka = [H+][CH3COO-]/[CH3COOH]

Ka = x*x/(c-x)

Assuming x can be ignored as compared to c

So, above expression becomes

Ka = x*x/(c)

so, x = sqrt (Ka*c)

x = sqrt ((1.738*10^-5)*4*10^-2) = 8.337*10^-4

since x is comparable c, our assumption is not correct

we need to solve this using Quadratic equation

Ka = x*x/(c-x)

1.738*10^-5 = x^2/(4*10^-2-x)

6.951*10^-7 - 1.738*10^-5 *x = x^2

x^2 + 1.738*10^-5 *x-6.951*10^-7 = 0

This is quadratic equation (ax^2+bx+c=0)

a = 1

b = 1.738*10^-5

c = -6.951*10^-7

Roots can be found by

x = {-b + sqrt(b^2-4*a*c)}/2a

x = {-b - sqrt(b^2-4*a*c)}/2a

b^2-4*a*c = 2.781*10^-6

roots are :

x = 8.251*10^-4 and x = -8.425*10^-4

since x can't be negative, the possible value of x is

x = 8.251*10^-4

So, [H+] = x = 8.251*10^-4 M

use:

pH = -log [H+]

= -log (8.251*10^-4)

= 3.08

Answer: option 5